Gordon Chemicals Company acquires a delivery truck at a cost of $31,000 on January 1, 2017. The truck is expected to have a salvage value of $4,000 at the end of its 4-year useful life. Compute annual depreciation for the first and second years under the declining-balance method.
Answer #1

Depreciation rate as per straight line method=100/4=25%/year

Hence depreciation as per double decline=2*Depreciation rate as per straight line method*Beginning value of each period

Year Beginning value Depreciation Ending value
1 31000 (2*25%*31000)=$15500 (31000-15500)=15500
2 15500 (2*25%*15500)=$7750.